Find the side of a square with an area of 625 ft.^2 by using a quadratic equation. I know that to find the area of a square you do length times width and all sides of a square are equal so I found the square root of 625, which was 25. But I am not sure on how to solve this problem using a quadratic equation. Thanks!!
Answer by stanbon(75887)

Let each side of the square be "x".
EQUATION:
Area = x^2 = 625
x^2 = 625
x^2-625 = 0
(x-25)(x+25)= 0
Positive answer:
x = 25
=============
Cheers,
Stan H.